### v2

# Unexplainability of Artificial Intelligence Judgments in Kant’s Perspective

Abstract
Kant’s Critique of Pure Reason, a major contribution to the history of epistemology, proposes a table of categories to elucidate the structure of the a priori principle of human judgment. The technology of artificial intelligence (AI), based on functionalism, claims to simulate or replicate human judgment. To assess this claim, it is necessary to study whether AI judgment possesses the characteristics of human judgment. This paper argues that AI judgments exhibit a form that cannot be understood in terms of the characteristics of human judgments according to Kant. Because the characteristics of judgment overlap, we can call this AI’s uncertainty. Then, I show that concepts without physical intuitions are not easy to explain when their functions are shown through vision. Finally, I illustrate that even if AI makes sentences through subject and predicate in natural language, which are components of judgment, it is difficult to determine whether AI understands the concepts to the level humans can accept. This shows that it is questionable whether the explanation through natural language is reliable.
Introduction
Philosophy has explored the concept of human intelligence for thousands of years, and the study of intelligence is a traditional philosophical topic. Logic attempts to explain the structure of knowledge, and epistemology seeks to understand how knowledge is obtained. In computer science, the powerful learning capabilities of deep learning (LeCun, Bengio, and Hinton 2015) have attracted attention. In particular, artificial neural networks (ANNs) (LeCun, Bengio, and Hinton 2015) are commonly used to mimic human intelligence. However, the information that has been produced with respect to the type of judgment used by artificial intelligence (AI) remains insufficient. Indeed, the nature of the judgment made by AI is often considered a black box (Adadi and Berrada 2018). Philosophy, with its tools to explain and analyze intelligence, should investigate the mimicry of the intelligence of ANNs.
First, I examine this topic broadly using Kant’s four categories of judgment to show that unexplainability can occur in all of them. Then, explainable AI technology is introduced. I then discuss the SoftMax function (Sharma, Sharma, and Athaiya 2017) that can force AI judgments to make possibility judgments. Next, the limitations of vision-based explainable AI technology are examined based on the fact that concepts whose functions are revealed in vision but have no physical intuitions are difficult to explain. The self and self-consciousness are taken as examples. Finally, I show that even by using natural language processing technology that reveals subjects and predicates for judgments a machine makes, it is difficult to determine whether a machine understands the judgments it makes at a level that a human would accept. Therefore, this shows that explanation through natural language processing is not reliable.
Background
Kant
(Kant 1855), which translated Kant’s book into English, is being used in this work because Kant’s Critique of Pure Reason was not originally written in English. This paper investigates AI judgments. For this, it is necessary to understand what judgment is. Judgments are divided in philosophy into subjects and predicates. For example, in the sentence “Socrates is wise,” “Socrates” is the subject, and “is wise” is the predicate. Basic information about judgment can be obtained in Kant’s discussion of analytic and synthetic judgment (Kant 1855, p. 7).
In all judgments wherein the relation of a subject to the predicate is cogitated, (I mention affirmative judgments only here; the application to negative will be very easy,) this relation is possible in two different ways. Either the predicate B belongs to the subject A, as somewhat which is contained (though covertly) in the conception A; or the predicate B lies completely out of the conception A, although it stands in connection with it. In the first instance, I term the judgment analytical, in the second, synthetical. Analytical judgments (affirmative) are therefore those in which the connection of the predicate with the subject is cogitated through identity; those in which this connection is cogitated without identity, are called synthetical judgments. The former may be called explicative, the latter augmentative judgments; because the former add in the predicate nothing to the conception of the subject, but only analyse it into its constituent conceptions, which were thought already in the subject, although in a confused manner; the latter add to our conceptions of the subject a predicate which was not contained in it, and which no analysis could ever have discovered therein. For example, when I say, “All bodies are extended,” this is an analytical judgment. For I need not go beyond the conception of body in order to find extension connected with it, but merely analyse the conception, that is, become conscious of the manifold properties which I think in that conception, in order to discover this predicate in it: it is therefore an analytical judgment. On the other hand, when I say, “All bodies are heavy,” the predicate is something totally different from that which I think in the mere conception of a body. By the addition of such a predicate therefore, it becomes a synthetical judgment. Judgments of experience, as such, are always synthetical.
In his exploration of the possibility of synthetic a priori judgments, Kant intricately maps the capabilities of intelligence, and he illustrates and defines the forms of judgment (Kant 1855, p. 58).
If we abstract all the content of a judgment, and consider only the intellectual form thereof, we find that the function of thought in a judgment can be brought under four heads, of which each contains three momenta.
Based on this claim, he presents Table 1 describing the features of judgments. Only through these can our intellect understand and think.
Quantity Universal Particular Singular
Quality Affirmative Negative Infinite
Relation Categorical Hypothetical Disjunctive
Modality Problematical Assertorical Apodeictical
Table 1: Logical function of the understanding in judgments (Kant 1855, p. 58).
Table 1 can make us assign each judgment to a specific category. For example, the proposition “All the values of 36 in pixels made by an infrared camera might indicate an actual human,” relates to universal, affirmative, categorical, and problematical.
Kant’s categories could be applied to natural language processing (Chowdhary and Chowdhary 2020) in computer science. However, this has limitations due to the limitations of Kant’s table, which is primarily designed for use with propositions having truth values. It seems to be applicable to the field of image captioning (Hossain et al. 2019) with truth values. However, in general, orders and exclamations lack truth value in natural language. There are also instances in which propositions are hidden through irony. Consequently, natural language processing that is based on logic cannot possess the same capabilities as humans due to this inherent limitation.
Artificial intelligence
A. M. Turing provided a striking test to assess the question of whether machines can think which is now known as the Turing test (French 2000; Turing 2009). According to this, if a machine provides responses that are indistinguishable from the way that a human would respond, then the two should not be differentiated. This aligns with functionalism (Levin 2004) in the philosophy of mind.
What makes something a thought, desire, pain (or any other type of mental state) depends not on its internal constitution, but solely on its function, or the role it plays, in the cognitive system of which it is a part.
Computer scientists generally and unwittingly adopt functionalist perspectives on the human mind. Many of these scholars contend that when computers perform human-like functions, we can recognize them as AI.
As a prominent AI algorithm, ANNs incorporate interconnected artificial neurons. Generally, ANNs are trained to generate correct output values that correspond to certain given input values. The significance of ANNs as a representative algorithm is evident in studies demonstrating their ability to efficiently train approximate solutions for all functions (Csáji et al. 2001; Hornik, Stinchcombe, and White 1989; Schäfer and Zimmermann 2006; Zhou 2020). This implies that ANNs can find all functions, and therefore, it can be assumed that all functions can be computed by using ANNs.
I would like to emphasize that the process of forward propagation, through which artificial intelligence makes judgments, is completely mathematically explained. However, what judgments are made is not explained like human judgment. During their training, all of the error contributions generated by neurons are calculated using a backpropagation algorithm (LeCun, Bengio, and Hinton 2015), which enables the ANN to learn by minimizing errors. This backpropagation algorithm provides information in a readable form using numbers. However, this form cannot be considered to be the equivalent of a human judgment. We know why AI makes judgments in numerical format, but we do not know expressions resembling human judgment. The field that aims to interpret AI judgments like human judgments is called explainable AI (Adadi and Berrada 2018).
Unexplainability of AI based on Kant’s epistemology
When AI makes a judgment, it produces numerical outputs. For instance, in image classification, let us say that if an image contains a face, the AI outputs 1; otherwise, it outputs 0. However, the determination of whether the judgment corresponds to the form of “Universal,” “Particular,” or “Singular” in Table 1 is challenging. For instance, when an ANN learns to classify a human face, people may require additional information to distinguish among statements such as “This image has a human face,” “Some of these types of images have a human face,” and “All these types of images have a human face.” Human beings’ expressions of their judgments are made in their exact form with natural language. However, AI outputs are represented as numbers. Therefore, when AI makes a specific judgment, we cannot classify its form in quantity in Table 1.
In logic, only “Affirmative” (A is B) and “Negative” (A is not B) are deemed necessary. However, Kant distinguishes “Infinite” from “Negative” (A is (not B)) (Kant 1855, p. 59).
If I say of the soul, “It is not mortal,”—by this negative judgment I should at least ward off error. Now, by the proposition, “The soul is not mortal,” I have, in respect of the logical form, really affirmed, inasmuch as I thereby place the soul in the unlimited sphere of immortal beings. Now, because, of the whole sphere of possible existences, the mortal occupies one part, and the immortal the other, neither more nor less is affirmed by the proposition, than that the soul is one among the infinite multitude of things which remain over, when I take away the whole mortal part. But by this proceeding we accomplish only this much, that the infinite sphere of all possible existences is in so far limited, that the mortal is excluded from it, and the soul is placed in the remaining part of the extent of this sphere. But this part remains, notwithstanding this exception, infinite, and more and more parts may be taken away from the whole sphere, without in the slightest degree thereby augmenting or affirmatively determining our conception of the soul.
It is not clear whether ANN outputs correspond to forms of “Affirmative,” “Negative,” or “Infinite.” For example, let us take again AI that is trained to output 1 when an image contains a human face and 0 when it does not. The judgments it makes during training with a value of 0 could be interpreted as “This image has a nonhuman face” or “This image does not have a human face.”
Let me examine several cases in detail. While “This image has a human face,” “This image does not have a human face,” and “This image has a nonhuman face” are all plausible interpretations, some are ruled out as a result of the characteristics of the task. First, we can imagine that the AI only uses the “Affirmative” form of a judgment when it outputs 1 or 0. When it outputs 1, it states, “This image has a human face.” However, when it outputs 0, no predicate can be affirmed, except for the affirmation of the negative predicate itself. Hence, we conclude that ANN utilizes a combination of the “Affirmative” and “Infinite” forms of judgment. Second, we can imagine that the AI makes judgments that only take the “Negative” form. Here, an output of 0 would mean that “This image does not have a human face,” and an output of 1 would not produce any corresponding predicate, with the exception of the “Affirmative” form. Hence, judgments are only possible involving a combination of “Affirmative” and “Negative” judgments. The use of only the “Infinite” form does not provide a solution. Additionally, with the exception of the two combinations of judgments determined above, a combination of the “Infinite” and “Negative” judgments would also seem impossible. If the ANN utilizes a combination of the “Affirmative” and “Infinite” forms of judgment, the question then arises: Is this type of learning acceptable? The ANN could be affirming that something is a nonhuman face but lacks further specification.
From a logical point of view, the “Infinite” form is identical to the “Affirmative” form, and previous studies have found that ANNs perform their tasks effectively: They should output true values where we intended the output to be true and false where we intend the output to be false. However, due to the differences in content according to the quality form of the judgments, an ANN may decide not to make judgments in the way that we intend. For instance, it could learn to make judgments on an entirely different basis from humans. An ANN may process this in a mixed way, utilizing both “Negative” and “Infinite” forms.
Considering Kant’s relation of judgment, this case presents even more difficult challenges than before. As summarized in Table 1, “Categorical” represents judgments on two concepts, “Hypothetical” indicates the relation between the two judgments, and “Disjunctive” signifies the relation between multiple judgments. For example, let’s consider a model that classifies objects as either a car, a dog, or a cat. It’s not always clear whether the model’s judgment is based on a single proposition, a causal relationship, or multiple propositions. The judgment could simply be “This image has a dog.” It could also be “This image has a dog because the image has a specific color.” Also, it could be “It is not a car and cat. I don’t know what a dog is. But it has to be a dog anyway. Therefore, there must be a dog.” In this last case, AI can make an accurate judgment without even possessing a full understanding of the concept of a dog. If this kind of learning occurs, it suggests that the learning process doesn’t align with our intention. Considering that AI is typically trained to classify objects within specific predefined classes, this phenomenon is quite common. Even when employing various technologies to elucidate decisions by highlighting the dog’s location in an image, it is uncertain whether AI comprehends the concept of a dog. This uncertainty arises because the AI might have made this determination based on distinguishing it from cats and cars.
[…]
$\vee¬(if ¬A, then ¬B)$
if (if A, then B), then ¬(if ¬A, then ¬B)
Table 2: Solution for the XOR problem.
Figure 1: Multilayer perceptrons solving the XOR problem.
Figure 1 shows that, although the value of true is defined as 1 and false as 0, there is no way to know whether AI has solved a logic problem or a classification problem. ANNs can solve the problem as a logical problem, as shown in Table 2, or as a problem of classification. Thus, we observe that unexplainability can exist in any of the four categories of judgment. Inspired by the uncertainty principle in quantum mechanics (Busch, Heinonen, and Lahti 2007), we can call this AI’s uncertainty. In this case, all forms of judgment overlap.
In the field of explainable AI research, various technologies are being studied (Adadi and Berrada 2018; Selvaraju et al. 2017). I believe that by developing technologies that provide insights into how AI judgments are made, we can achieve some level of explainability. Computer scientists have begun revealing which part of an image AI focuses on when making judgments. These tools assist us in inferring the subject and predicate. However, neither a subject nor a predicate is explicitly stated.
Among the four categories that are shown in Table 1, modality seems easier to put into explainable form than any of the others. This is because of the SoftMax function (Sharma, Sharma, and Athaiya 2017), which is applied to the final layer, ensuring that the cumulative sum of all options equals 1. This makes outputs satisfy the definition of probability. For example, when looking at an image and determining whether there is a dog or a cat, the SoftMax function makes it possible to indicate a 63% probability that there is a dog and a 37% probability that there is a cat. According to Kant, the modality form has the following characteristics (Kant 1855, pp. 60–61).
The modality of judgments is a quite peculiar function, with this distinguishing characteristic, that it contributes nothing to the content of a judgment (for besides quantity, quality, and relation, there is nothing more that constitutes the content of a judgment), but concerns itself only with the value of linking verb in relation to thought in general.
Because it does not contribute to the content of the judgment, it allows us to explain it in the “Problematical” form even if we do not know what judgment the machine makes. Therefore, even before the advent of explainable AI, it was possible to obtain judgments in the form of “Problematical” only through the SoftMax function. What we want to know through explainable AI may be the characteristics of the entire judgment shown in Table 1, but another way is to artificially consider only the content of the judgment and intentionally fix the modality of the judgment. In other words, it is not that we will see what kind of judgment was made, but that we will treat it as a possibility judgment no matter what judgment was made. People might claim that we can transform some judgments into possibility judgments. For example, if something is said to be inevitably and logically true, the probability is 100%. Can “actually exist,” “necessary,” and “possible” be the same? Treating all judgments as possibility judgments seems problematic.
Limitation of vision-based explainable AI
We may encounter a range of things that cannot be explained when dealing with concepts that have no visual intuitions but whose functions are revealed in vision. I address this topic by beginning with an experiment conducted by (Seo and Lee 2021). To predict the speed of a ball, it is crucial to observe its movement. However, if the camera that is filming it also begins to move in the same manner the ball does, the ball will always appear stationary in the camera’s images. Hence, the position of the camera must also be taken into account. Coordinate information, such as those provided by the global positioning system (El-Rabbany 2002) can be utilized in this regard. Moreover, the camera’s position can be estimated by using fixed background information. If the background screen moves, it means that the camera moves. If it moves only to a minimal degree, the camera moves only minimally, Figure 2 provides an image of this scenario.
Figure 2: A ball, camera, and colorful boxes fixed in the background.
Performance decreased when boxes were eliminated (Seo and Lee 2021). This indicates that the AI used boxes to predict the speed. From this, I conducted additional experiments. Specifically, I seek to demonstrate that the AI utilizes the boxes for information from boxes by employing Grad-CAM (Selvaraju et al. 2017). Grad-CAM focuses on a classification problem, so it cannot be used for the task that predicts the speed itself. Therefore, it is trained to classify class 1 for high speeds based on the average of final speeds in the training set and class 2 for low speeds. When AI predicts speeds higher than the average, I examine where it focuses on its attention.
I employ a combination of a convolutional neural network (CNN) (Lawrence et al. 1997), which processes two-dimensional spatial information (e.g., photos), and a long short-term memory (LSTM) (Hochreiter and Schmidhuber 1997) which processes data obtained along the time axis. This is called a long-term recurrent convolutional network (LRCN) (Donahue et al. 2015). The training data comprise 200 videos, each containing 20 images captured at 0.2-second intervals. The trained model consists of 4 CNN layers, 2 LSTMs, and 1 fully connected layer (LeCun, Bengio, and Hinton 2015). If the values in parentheses represent the input channel, the output channel, the square kernel size, and the stride in order, the parameters for the CNN layers are (3,15,5,2), (15,13,5,2), (13,7,5,2), and (7,3,5,2). The LSTM has two layers and outputs 40 nodes. Finally, the LRCN outputs the speed in the last fully connected layer. The learning rate is set to 1e-4, and Adam (Kingma and Ba 2014) is used. Cross entropy loss function (Wang et al. 2020a) is employed, and the learning epoch is set to 2000. Figure 3 presents the result. I use the activation map of the CNN layer that is first used in the forwarding process, which processes input values.
Figure 3: Where does AI focus?
Figure 3 displays the areas where AI focuses on. The area marked in white is the area that AI focuses on. These white areas are near the balls and the boxes. This indicates that the speed of the ball is predicted in relation to the information from the balls and boxes, as discovered by (Seo and Lee 2021). It is intriguing that even though I only provided the correct classes for the speed of the ball and the video, the AI began to give its attention to the balls and boxes, indirectly utilizing the boxes for self-location.
We could imagine the machine thinking as follows: While tracking the position of the ball and the box, considering that my location has moved by this much, the speed is likely to be at this level. Here, we ask the crucial question: Does the fact that a machine seeks self-location mean that it understands the concept of self? Humans address the problem using self-consciousness. Do machines also have such concepts, based on functionalism? Problems arise from the following. The concepts of dogs and cats have physical intuitions. However, it seems that the concept of self has no such accompanying intuitions. Let us investigate the concepts relating to humans, animals, and AI.
We can think that the structure of human cognition is the same as each other, assuming no illness is present. The cognitive structure itself is completely given and becomes fully experienced. Research seeks to elucidate how cognition functions, leading to investigations of epistemology and logic. In epistemology, Kant argues for the division of self-consciousness into transcendental and empirical aspects. This perspective and other questions persist in philosophy and in the empirical sciences. First of all, several attempts have been made to scientifically examine our self-consciousness. Bodily self-consciousness has been explored (Blanke 2012; Blanke, Slater, and Serino 2015; Ionta et al. 2011; Lenggenhager, Mouthon, and Blanke 2009; Lenggenhager et al. 2007; Serino et al. 2013). In particular, (Blanke 2012) includes attempts to determine the neural mechanisms that underlie self-identification, self-location, and the first-person perspective. The question of whether individuals with autism possess the concept of self and how it can be enhanced has been investigated (Duff and Flattery Jr 2014; Huang et al. 2017; Lind and Bowler 2009; Lombardo et al. 2007; Mitchell and O’Keefe 2008). On the other hand, Kant discussed self-consciousness in the following paragraph (Kant 1855, pp. 81–83).
The I think must accompany all my representations, for otherwise something would be represented in me which could not be thought; in other words, the representation would either be impossible, or at least be, in relation to me, nothing. […] For the manifold representations which are given in an intuition would not all of them be my representations, if they did not all belong to one self-consciousness, that is, as my representations (even although I am not conscious of them as such), they must conform to the condition under which alone they can exist together in a common self-consciousness, because otherwise they would not all without exception belong to me. […] For example, this universal identity of the apperception of the manifold given in intuition, contains a synthesis of representations, and is possible only by means of the consciousness of this synthesis. For the empirical consciousness which accompanies different representations is in itself fragmentary and disunited, and without relation to the identity of the subject. This relation, then does not exist because I accompany every representation with consciousness, but because I join one representation to another, and am conscious of the synthesis of them. Consequently, only because I can connect a variety of given representations in one consciousness, is it possible that I can represent to myself the identity of consciousness in these representations; in other words, the analytical unity of apperception is possible only under the presupposition of a synthetical unity. The thought, “These representations given in intuition, belong all of them to me,” is accordingly just the same as, “I unite them in one self-consciousness, or can at least so unite them;” and although this thought is not itself the consciousness of the synthesis of representations, it presupposes the possibility of it; that is to say, for the reason alone, that I can comprehend the variety of my representations in one consciousness, do I call them my representations, for otherwise I must have as many-coloured and various a self as are the representations of which I am conscious.
When a human solves the task given to the AI, it is necessary that all of the intuitions of all images coming in at every moment belong to a single self-consciousness. Because, if I borrow a phrase from Kant, self-consciousness is conscious of the synthesis of representations. Therefore, self-consciousness is needed to solve the task. And the concept of self is needed for considering self-location.
Animals cannot directly share their experiences with us, so we are in need of evidence to clarify that they are aware of the concepts of self and self-consciousness. Kant also grappled with this difficulty (McLear 2011). Animals seem to have a concept of self. The empirical investigations that have researched this concept in animals have primarily focused on self-location. For instance, place cells, which are found in the hippocampus of mice are known as cells for self-location (Barry and Burgess 2014; Chen et al. 2019; Danjo, Toyoizumi, and Fujisawa 2018; Knierim 2015; Nadel and Eichenbaum 1999). Place cells are activated with respect to the mouse’s location. Specific neurons are activated in specific spatial positions, and these change as the mouse moves. Also, Mice acquire self-location through background information (Chen et al. 2019). The 2014 Nobel Prize in Physiology or Medicine was awarded for the discoveries of the cell. I want to emphasize that animals’ ability to understand self-location is not a mere claim but a well-established fact. These studies might be pieces of evidence that animals process temporal and spatial information, implying the existence of the concepts of self and self-consciousness.
The experiment reported in Figure 3 could help understand the acquisition of the concept of self and self-consciousness in humans and animals. This concept appears to have evolved from the predator–prey relationship. In Figure 3, we can regard the ball as either the prey or the predator and the camera as the opposite in each case. This situation could aid the development of the concepts of self and self-consciousness. Some studies (Bard et al. 2006; Gallup Jr, Anderson, and Shillito 2002; Haikonen 2007) have shown that if an animal can recognize itself in a mirror, it has the concept of self-consciousness. However, mirrors do not exist in nature except still water surfaces. I believe that the concepts of self and self-consciousness can be derived from the relationship between predator and prey.
What sets AI apart from humans and animals is that we can create its cognition structure to be similar to our own, as it is a technology we create, not a natural science that must be verified empirically. This leads us to interpret that AI has a cognitive structure that resembles ours, even if only partially, provided that sufficient conditions are provided. These characteristics can lead to attaching the word “Artificial” to an existing concept. Explainable AI involves a long journey that involves identifying the parts of AI that align with human cognition and striving to make it similar to ours. However, AI researchers are not obligated to always make AI explainable.
Based on this generous attitude, previous investigations (Bringsjord et al. 2015; Chella, Frixione, and Gaglio 2008; Zhang and Zhou 2013) have argued that machines have self-consciousness. However, we cannot have any confidence in the concepts that do not have visual intuitions. These concepts could be explainable in other ways, but they cannot be interpreted in a vision-based explainable way. We can only confidently conclude that there is a relationship between these and other parts. Even if vision-based explainable technology develops, it cannot explain anything about concepts that appear in vision but do not have physical intuitions.
We can emphasize that AI is created to perform specific tasks. In natural language processing, the chat of AI is an important topic. When a certain sentence is inputted, the AI outputs a response (Deng et al. 2019; Zhou et al. 2018). Recent technologies have displayed chatbots that can pass the Turing test. Nevertheless, we do not assert that chatbots have self-consciousness. They are just set up to behave that way. Using human chat data, AI is trained to output appropriate sentences. As another example, a generative adversarial network consists of a discriminator and a generator (Creswell et al. 2018; Goodfellow et al. 2014). Here, the generator produces random fake data, and the discriminator is responsible for distinguishing between the real and the fake data. Making them compete with each other, we train the generator to produce data that are as similar to real data as possible. For another example, reinforcement learning studies how an agent adapts to the environment through changing actions and states after dividing a system into an agent and environment (Mazyavkina et al. 2021; Sutton and Barto 2018). Both cases involve humans dividing two groups in advance and making them interact. I doubt whether these computer programs acquire the concept of the self in this process. Even without this concept, they can address problems because we specifically ask them to do this. There is no inherent reason why they must possess these concepts.
Vision, natural language, and definition in Kant’s perspective
When explaining AI judgment, in addition to highlighting pixels, it could be helpful to have a textual output. But we could have doubts. To make a judgment, AI must understand the concepts used in a judgment. Does the machine truly understand the concepts expressed in natural language? In other words, does AI accurately understand the definitions of all the concepts in sentences? Kant said the definition as follows (Kant 1855, pp. 443–444).
A definition is, as the term itself indicates, the representation, upon primary grounds, of the complete conception of a thing within its own limits. Accordingly, an empirical conception cannot be defined, it can only be explained. For, as there are in such a conception only a certain number of marks or signs, which denote a certain class of sensuous objects, we can never be sure that we do not cogitate under the word which indicates the same object, at one time a greater, at another a smaller number of signs. Thus, one person may cogitate in his conception of gold, in addition to its properties of weight, colour, malleability, that of resisting rust, while another person may be ignorant of this quality. We employ certain signs only so long as we require them for the sake of distinction; new observations abstract some and add new ones, so that an empirical conception never remains within permanent limits. […] In the second place, no a priori conception, such as those of substance, cause, right, fitness, and so on, can be defined. For I can never be sure, that the clear representation of a given conception (which is given in a confused state) has been fully developed, until I know that the representation is adequate with its object. But, inasmuch as the conception, as it is presented to the mind, may contain a number of obscure representations, which we do not observe in our analysis, although we employ them in our application of the conception, I can never be sure that my analysis is complete, while examples may make this probable, although they can never demonstrate the fact. […] Consequently, the science of mathematics alone possesses definitions. For the object here thought is presented a priori in intuition; and thus it can never contain more or less than the conception, because the conception of the object has been given by the definition—and primarily, that is, without deriving the definition from any other source.
Every year, hundreds of papers are published in numerous journals to report new findings on neural networks in biology. Although we do not have a complete understanding of neural networks, we have made the definition of neural networks from our limited knowledge. Can such a definition be complete? Nevertheless, most people admit that ANNs have been created that resemble neural networks. In Kant’s context, although it may seem confusing, there exists a sufficiently agreed definition. We can infer that the definition of an empirical concept relies on experience and appears to be closely intertwined with consensus and language utilization. The idea that meaning is determined by the way language is utilized is well expressed in Wittgenstein’s philosophy of language (Travis 1989) and plays an important role in the philosophy of language as well. It is worth noting that natural language processing is also designed with the view that the meaning of a word is determined by the existence of surrounding words (Devlin et al. 2018; Le and Mikolov 2014; Mikolov et al. 2013; Vaswani et al. 2017).
Even if a machine cannot strictly know concepts precisely, if it is true that AI vaguely understands the concepts, then AI understanding them should be admitted. However, it is hard to determine whether AI understands the concepts in the sentences made by AI to the degree that would be generally accepted by humans. For example, it does not seem easy to determine whether AI understands the definitions and meanings of concepts such as “Is,” “Have,” and “This” to a level that humans can accept. In situations where it is questionable whether AI understands each concept, we cannot trust AI judgments shown in sentence form.
Conclusion
If AI is intended to imitate human judgment, it should exhibit the categories of human judgment. According to Kant, four such categories exist. I analyzed AI judgments within each. In the result, it is shown that AI judgments are ambiguous. Because the characteristics of judgments overlap, we can call this AI’s uncertainty. Furthermore, this paper demonstrates that concepts lacking physical intuitions are challenging to elucidate when they are visually represented through their functions. Lastly, I show that even when utilizing natural language processing technology for expression through subject and predicate, which are fundamental components of judgment, it is hard to determine whether AI genuinely comprehends the concepts it employs to a degree acceptable to humans. Therefore, we cannot trust AI judgments.
We have striven to understand AI judgments within the confines of our human intellect. Can AI judgments possess additional characteristics beyond those of human judgments? Even if such judgments can exist, it appears that we lack the means to fully comprehend them. All problems stem from our acknowledgment that AI makes judgments.